OpenStudy (anonymous):

How many arrangements of 3 digits can be formed from the digits 0 through 9?

OpenStudy (anonymous):

729??

OpenStudy (anonymous):

Let me check..

OpenStudy (lgbasallote):

hmm doesnt sound right...

OpenStudy (anonymous):

9*10*10 = ??

OpenStudy (anonymous):

900

OpenStudy (lgbasallote):

0 to 9 has 10 digits. you take 3 from that...so it will be \[10C3 \implies \frac{10!}{3!(10-3)!}\] right?

OpenStudy (anonymous):

Arrangements @lgbasallote not selection only..

OpenStudy (lgbasallote):

so permutation?

OpenStudy (anonymous):

Yes permutation..

OpenStudy (lgbasallote):

9*10*10 doesnt sound like permutation...

OpenStudy (anonymous):

720??

OpenStudy (anonymous):

See I have simply applied the concept..

OpenStudy (lgbasallote):

\[10P3 \implies \frac{10!}{(10-3)!}\] right? @waterineyes

OpenStudy (anonymous):

Solve that..

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

But you cannot fill first place with 0..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

See: Suppose three digit number is : ABC Here A you can fill with : 1 to 9. B can be filled with 0 to 9 Similarly C can be filled with 0 to 9 digits: So how may ways: 9*10*10 = 900..

OpenStudy (anonymous):

These are the numbers with repetition..

OpenStudy (lgbasallote):

but it didnt specifically say that you cant use 0 as first digit..

OpenStudy (anonymous):

Without repetition: 9*9*8 = ??

OpenStudy (anonymous):

If you use 0 as first digit then it will become two digit number and not three digit. It is understood..

OpenStudy (lgbasallote):

not really..no...

OpenStudy (anonymous):

Yes I am sure..

OpenStudy (anonymous):

Without repetition you will get three digit numbers = 576 And with repetition you will get : 900..